OpenStudy (anonymous):

What are all the subsets of the set? (-3, 6)

OpenStudy (kmeis002):

Do you mean {-3, 6}. Dont mean to be a stickler, but the difference in brackets really makes a difference in the answer.

OpenStudy (anonymous):

YES sorry... They are Brackets. I couldn't see them with the font on my computer.

OpenStudy (kmeis002):

Given a set A a subset of A is such that all elements in the subset are contained in A. Ex: Let A = {a, b,c d} then B = {a,b} is a subset because a contains a and b. So it you have A = {-3, -6} Then {-3} is a subset, {6} is a subset, and so is {-3,6} (even though its the same set, by definition a set is a subset of itself). Note, there is always one more subset, the empty set {}.

OpenStudy (anonymous):

Thank you so Much!!!! :D

OpenStudy (anonymous):

\[sub-sets of \left\{ 3,-6 \right\}=\emptyset,\left\{ 3 \right\},\left\{ -6 \right\},\left\{ 3,-6 \right\}\] \[no. of sub-sets=2^{no. of elements of the set.}\]